OpenStudy (anonymous):

the probability that a professional baseball player will get a hit is 1/3. calculate the exact probability that he will get at least 3 hits in 5 attempts

OpenStudy (anonymous):

at least 3 in 5 attempts means 3 or 4 or 5 and you have to calculate these separately do you know how to do it?

OpenStudy (anonymous):

exactly 5 is easiest, it is \((\frac{1}{3})^5\)

OpenStudy (anonymous):

exactly 4 is \(\binom{5}{4}(\frac{1}{3})^4(\frac{2}{3})\)

OpenStudy (anonymous):

since \(\binom{5}{4}=5\) this is \(5\times (\frac{1}{3})^4\times (\frac{2}{3})\)

OpenStudy (anonymous):

do you know how to compute exactly 3?

OpenStudy (anonymous):

is exactly 3 = \[\left(\begin{matrix}40 \\ 243\end{matrix}\right)\]

OpenStudy (zarkon):

you should write it as a fraction \frac{40}{243} \[\frac{40}{243}\]

OpenStudy (anonymous):

hee hee thx, kinda a noob ^^;

OpenStudy (anonymous):

exactly 3 is \(\binom{5}{3}(\frac{1}{3})^3(\frac{2}{3})^2\) yes, it is \(\frac{40}{243}\) add up to get your answer

OpenStudy (anonymous):

\[\frac{51}{243}\] hope that's right :3

OpenStudy (zarkon):

yep...reduces to \[\frac{17}{81}\]

OpenStudy (anonymous):

thanks for the help! :D